The brain volumes (cm3) of 20 brains have a mean of 1118.2 cm3 and a standard deviation of 129.3 cm3 . Use the given standard deviation and the range rule of thumb to identify the limits separating values that are significantly low or significantly high. For such​ data, would a brain volume of 1326.8 cm3 be significantly​ high?

Significantly low values are ____ cm3 or lower.

Significantly high values are ____ cm3 or higher.
Answer #1

Significantly low values = Values below 2 standard deviation = (1118.2 - 2*129.3) cm3 or lower = 859.6 cm3 or lower

Significantly high values = Values above 2 standard deviation = (1118.2 + 2*129.3) cm3 or higher = 1376.8 cm3 or higher
Answer #2

The brain volumes

()

of 20 brains have a mean of

and a standard deviation of

.

Use the given standard deviation and the range rule of thumb to identify the limits separating values that are significantly low or significantly high. For such data, would a brain volume of

be significantly high?